OpenStudy (anonymous):

If f(x) = 4x + 1 and g(x) = x2 – 2x + 1, what is f(g(x)) ?

OpenStudy (anonymous):

\[f \left( g \left( x \right) \right)= 4g \left( x \right)+1 = 4\left( x ^{2}-2x+1 \right)+1= 4x ^{2}-8x +5\]